# MindShift: Leveraging Large Language Models for Mental-States-Based Problematic Smartphone Use Intervention

### v1

@@ 4.3. When to Intervene: Intervention Trigger Mechanism
As in Figure 2, each mental state allows for multiple persuasion strategies. Therefore, we devise a simple procedure. After determining the mental state and narrowing down the specific strategies, we first pick one strategy and generate an intervention message. Then, we loop over other different strategies and show new strategy messages every two minutes until the users leave the app. After looping over all appropriate strategies under this mental state, the intervention will stop. Users only need to report (i.e., Step 2 and 3 in Figure 5) the first time they open a specific app after each screen unlock. This design is based on two considerations: 1. We assume that the user’s initial opening usually reflects their subsequent usage intentions. 2. Requiring users to report multiple times when switching between apps can lead to a poor user experience.
@@ 5.1. Client-Side Implementation
The client-side is an Android app that implements all the features of our design. We use accessibility services to detect the opening and closing of apps on the phone. The client is also responsible for collecting and uploading data to the server, including the screen’s off and unlock status, the name, opening and closing times of the application, and location data obtained through the Amap API (Amap, 2023). To prevent the accidental killing of the accessibility service and ensure compliance, we also add a background service in the app to check the status of the accessibility service every 5 minutes. If it detects that the accessibility service is terminated, the client will notify the server to send an email to a researcher. The researcher can remind the user to turn on the accessibility service again to ensure the integrity of the data.
@@ 5.2. Server-Side Implementation
The server side is responsible for generating persuasive content tailored to the user’s context through four key tasks. By performing these tasks, the server side ensures that the intervention is personalized, contextually relevant, and delivered in a timely manner.
(1) User Data Computation: The server first computes the user data that needs to be used in the word slots from the app data uploaded by the client, including the total habitual usage time of the phone and the time of the last habitual opening.
(2) Habit Selection: Next, the server selects a habit mostly matched with the current user’s mental state, location (i.e., the specific building), and time (i.e., the hour of the day) from users’ initialization (i.e., Step 1 in Figure 5). To create a strong association between habit and context, as long as users do not thumb down or modify it, the same habit will be recommended in the same context. More details can be seen in Step 4 of 4.2.
(3) Strategy Counterbalance: To ensure that the number of times each strategy appears in each mental state is roughly the same, the server then counterbalances the strategy order in the prompt mentioned in Section 4.2.
(4) Content Generation: After obtaining the user contexts, habits, and persuasion strategy, the server uses the OpenAI GPT-3.5 API to generate appropriate intervention content. We adopted a streaming, character-by-character generation approach, allowing the persuasive content to start being displayed within 2 seconds.
@@ 6. Field Experiment
To evaluate the effectiveness of MindShift, we conducted a 5-week field experiment. In this section, we introduce experimental design, participant recruitment, and experiment procedure.
6.1. Baseline Intervention Methods
As MindShift is one of the first persuasion intervention systems leveraging an LLM to generate dynamic persuasion content, there are no comparable systems other than the traditional persuasion techniques. We compare MindShift against a persuasive reminder baseline (also as a pop-up window), one of the most commonly adopted intervention methods in commercial apps (Inc, 2023; Apple, 2021). To ensure the fairness of the comparison, the persuasion baseline is designed to be the same as the intent report step, as illustrated in Step 2 in Figure 5. We name this intervention as Baseline.
Moreover, in order to evaluate the effectiveness of the mental states and persuasion strategies proposed in Section 3, we further design a simplified version of MindShift by removing the mental states and persuasion strategies from the prompt design. The content generation will solely be based on users’ physical contexts and their app usage behavior. We name this group as MindShift-Simple. In total, we have three intervention methods to compare: Baseline, MindShift-Simple, and MindShift. Table 1 shows the comparison of these three intervention techniques.
Table 1. Comparison of Three Intervention Techniques
@@ 6.1. Baseline Intervention Methods
Intent Report LLM-powered Persuasion Mental-State-based Persuasion Strategies
@@ 6.2. Experiment Design
We adopt a within-subject design, with intervention techniques as independent variables (Baseline, MindShift-Simple, and MindShift). We design a 5-week field experiment. To measure users’ everyday phone usage behavior, the first week is set as the Baseline stage, followed by two weeks of one MindShift version and another two weeks of the other version. We counter-balance the order of MindShift-Simple and MindShift.
Our evaluation metrics include various aspects: (1) Intervention acceptance rate. We measure the percentage of times users accept the intervention and quit the blacklist app use when interventions are shown. (2) Intervention thumb-up rate. For MindShift-Simple, and MindShift that show persuasion content, users can provide feedback by pressing buttons (thumb-up or thumb-down, see Figure5 step 4 interface). (3) App usage behavior, which includes both app opening frequency and usage duration. (4) Subjective reports. At the beginning of the study and at the end of each intervention session, we distributed the Smartphone Addiction Scale (SAS) (Kwon et al., 2013) and the self-efficacy Scale (Schwarzer and Jerusalem, 1995). In addition, at the end of the study, we also conducted a brief semi-structured interview to gather user experiences and feedback on our intervention techniques. These metrics cover both the objective and subjective measures of the interventions.
@@ 6.3. Participants
We sent out recruitment material on social media platforms. We included a screening survey aiming to identify potential participants who showed signs of smartphone addiction and the willing to reduce their smartphone use. Specifically, besides basic demographics, we included four questions selected from the SAS and self-efficacy questionnaires. We also included questions about the willingness to reduce smartphone use, the extent of habitual phone use, future plans for the next five weeks, and a screenshot of phone usage time of the last week. We excluded users who (1) did not show signs of smartphone addiction (SAS sub-score ¡ 15), or (2) were unwilling to reduce smartphone use, or (3) had a total weekly usage time of less than 20 hours, or (4) those who had special plans such as long-term travel in the next five weeks (which may shift their phone usage patterns).
We received a total of 42 responses. We recruited 31 participants after the screening process. 6 participants voluntarily dropped out during the study. For the remaining participants, we divided them into groups according to counterbalanced intervention orders. We conducted a Kruskal-Wallis test analysis to ensure that groups had no significant difference in the SAS scores and self-efficacy scores. In the end, 25 participants completed the entire study (females=13, males=12, age=22±2 years). Among them, 17 participants were undergraduates, 5 were graduate students, and 3 were professionals.
@@ 6.4. Experiment Procedure
After all the participants signed the consent form, we held a 20-minute onboarding session online to familiarize participants with the research process and introduce the Android application. We explained in detail the meanings of each selection in the report interfaces (Step 2 and 3 in Figure 5). After the meeting, participants filled out the first SAS and self-efficacy questionnaires. The app was then deployed for a 5-week field experiment.
Before users started using one of two versions of MindShift (MindShift and MindShift-Simple), we provided participants with a tutorial explaining the three mental states (i.e., boredom, stress, and inertia) that they need to report. This ensured that they had an accurate and consistent understanding of the mental states, which in turn improved the accuracy and reliability of the reported data. At the end of the experiment, participants received a compensation of \$50 for their time.
@@ 7. RESULTS
During the five-week study, we collect 50,815 minutes of restricted app usage duration, and 54,467 restricted app opening events (7539, 23,769, 21,994, and 835 for habitual use, instrumental use, relax, and quit). We analyze the quantitative data collected via app and the qualitative data collected via questionnaires and interviews.
@@ 7.1. Intervention Acceptance Rate
The effectiveness of a persuasion strategy is directly measured by the rate of successful prevention of user engagement with the targeted application (i.e., intervention acceptance rate). We compare the overall acceptance rate (Sec. 7.1.1), acceptance rate for generated persuasion content (Sec. 7.1.2), and thumb-up rate (Sec. 7.1.3). Moreover, we also conduct a detailed analysis of the acceptance rate across strategies (Sec. 7.1.4), mental states and activities (Sec. 7.1.5).
@@ 7.1.1. Overall Acceptance Rate
MindShift and MindShift-Simple increase the overall acceptance rate significantly and MindShift achieve best. We first investigate the overall acceptance rate during the three intervention sessions, the total number of quit times divided by the total number of app visits (excluding instrumental uses and relaxation). Figure 6 shows the comparison, MindShift (35.2±24.1%) has a higher acceptance than MindShift-Simple (30.5±23.6%) and Baseline (12.7±17.4%). These differences are statistically significant according to a Friedman test ($\chi^{2}$(2) = 13.69, p ¡ .01). Three post hoc Wilcoxon signed-rank tests, corrected with Holm’s sequential Bonferroni procedure, indicate that MindShift vs. Baseline (V = 52, p ¡ .01) and MindShift-Simple vs. Baseline (V = 38, p ¡ .001) are statistically significantly different, but that MindShift vs. MindShift-Simple is not significant (V = 129, n.s.).
@@ 7.1.2. Acceptance Rate for Generated Persuasion Content
MindShift has a significantly higher persuasion acceptance rate than MindShift-Simple. The overall acceptance rate includes two parts: exiting when reporting intention and exiting after seeing the persuasion content. To narrow down the comparison between MindShift and MindShift-Simple, we focus on the intervention acceptance rate during the persuasion stage, as the two intervention techniques’ differences lie only in the persuasive content. As shown in Figure 6(a), MindShift achieves higher persuasion acceptance (20.1±20.2%) than MindShift-Simple (12.0±15.0%) and a paired-samples t-test shows that MindShift was statistically significantly higher (t = -2.21, p¡0.05).
Moreover, as we introduce in Section 4.3, every persuasion intervention could consist of 1 to 4 rounds, depending on which mental state participants are in and which stage participants leave the app. Therefore, we further compare the persuasion acceptance rates between MindShift and MindShift-Simple across different rounds. The results show that MindShift outperforms MindShift-Simple at each round ($\Delta_{\textit{Round1}}$=6%, $\Delta_{\textit{Round2}}$=14.7%, $\Delta_{\textit{Round3}}$=10.2%, $\Delta_{\textit{Round4}}$=11%) as indicated in Figure 6(b). Moreover, we observe a decreasing trend of the acceptance rate of MindShift-Simple, while MindShift does not show such a trend. We conduct a paired-samples t-test between the two intervention techniques in each round. The results indicate that, except for the first round, rounds 2 (p¡0.05), 3 (p¡0.001), and 4 (p¡0.001) all exhibit that MindShift achieves significantly higher acceptance rate.

### v2

@@ 4.3. When to Intervene: Intervention Trigger Mechanism
We consider the user’s intent of use in the intervention trigger rule (Step 1 of Sec. 4.2). Specifically, interventions will be triggered when a user’s self-report intent is habitual use.
As in Figure 3, each mental state allows for multiple persuasion strategies. We devise a simple procedure. After determining the mental state and narrowing down the specific strategies, we first randomly sample one strategy and generate an intervention message. Then, we loop over other strategies and show new strategy messages every two minutes until the users leave the app. After looping over all appropriate strategies under this mental state, the intervention will stop. The usage duration is calculated based on the total usage time of a single blacklisted app during one unlocking session. The two-minute interval setting is derived from the statistical analysis in the WoZ study, where 90% of users spent less than 5 minutes on a blacklisted app in a single session. We thus set the interval between interventions as two minutes as a convenient delay to facilitate the exploration of different strategies, as further supported by the analysis in Sec. 7.2.2.
Users only need to report their habitual use and mental states once (i.e., Step 1 and 2 in Figure 5) when they open a specific app during each screen unlock session. This design is based on three considerations: (1) The initial mental state when opening an app is crucial, as it triggers habitual use (Sec. 3.2). (2) As reflected in our pilot study in Sec. 4.2, multiple reports during app switching can be annoying and negatively affect user experience. (3) Some previous studies suggest that users’ stress level tends to be retained even with coping techniques (Duvenage et al., 2020). We assume this also applies to other mental states, so most users’ mental state remains stable during one session (90% was less than 5 minutes), as further supported by the analysis in Sec.  7.2.1. We also discuss future potential ways to enhance accuracy in Sec. 8.3.2 & 8.4.
@@ 5.1. Client-Side Implementation
The client-side is an Android app that implements all the features of our design. We use accessibility services to detect the opening and closing of apps on the phone. The client is also responsible for collecting and uploading data to the server, including screen’s off and unlock status, application name, application opening and closing times, and location data obtained through the Amap API (Amap, 2023). To prevent the accidental killing of the accessibility service and ensure compliance, our app includes a background service checking the accessibility service status every 5 minutes. If it detects that the service is terminated, the app informs the server to email a researcher, who then reminds the user to reactivate the service, ensuring data integrity.
@@ 5.2. Server-Side Implementation
The server side is responsible for generating persuasive content through four key tasks, ensuring that the intervention is personalized, contextually relevant, and delivered in a timely manner.
(1) User Data Computation: The server processes user data from client-uploaded app data for use in word slots, including the phone’s total habitual use time and last habitual opening time.
(2) Habit Selection: Next, the server selects a habit mostly matched with the current user’s mental state, location (i.e., the specific building), and time (i.e., the hour of the day) from the users’ initialization (i.e., Step 1 in Figure 5). To reinforce the habit-context link, unless users thumb down or modify it, the same habit will be recommended in the same context. More details can be seen in Step 4 of 4.2.
(3) Strategy Counterbalance: To balance the frequency of each strategy across mental states, the server counterbalances strategy order in the prompt in Sec. 4.2.
(4) Content Generation: After obtaining the user contexts, habits, and persuasion strategies, the server uses the OpenAI GPT-3.5 API to generate persuasive content. We adopted a streaming, character-by-character generation approach, allowing the persuasive content to start being displayed within 2 seconds.
@@ 6. Field Experiment
To evaluate the effectiveness of MindShift, we conducted a 5-week field experiment. We introduce experimental design (Sec. 6.1 & 6.2), participant recruitment (Sec. 6.3), and experiment procedure (Sec. 6.4).
6.1. Baseline and MindShift-Simple Intervention Methods
As MindShift is one of the first persuasion intervention systems leveraging an LLM to generate dynamic persuasion content, there are no comparable systems other than the traditional persuasion techniques. We compared MindShift against a persuasive reminder baseline, one of the most commonly adopted intervention methods in commercial apps (Inc, 2023; Apple, 2021). To ensure the fairness of the comparison, the baseline is designed to be the same as the intent report step, as illustrated in Step 1 in Figure 5. Specifically, it only requires users to report the intent the first time they open the blacklist app after unlocking. It can also be used to collect the proportion of users’ initial intents, facilitating the analysis of the intervention effect of MindShift. We name this intervention as Baseline.
Moreover, in order to evaluate the effectiveness of the mental states and persuasion strategies proposed in Sec. 3, we further designed a simplified version, MindShift-Simple, by removing the mental states and persuasion strategies from the prompt design. Specifically, in the prompt design (Figure 4), we retained only ¡User Data¿ in the (2) Description of the current context and removed the (4) Description of the persuasion strategy. Meanwhile, the last sentence in ¡Output Format¿ was changed to generate four sentences at once. We kept other setups consistent, ensuring that both versions’ language features (such as both tone styles are humorous and caring) are as consistent as possible. Examples of content generation in MindShift-Simple are as shown in Table 2 under ‘With Context’ and ‘Without Strategy’.
In total, we have three intervention methods to compare: Baseline, MindShift-Simple, and MindShift. Table 3 shows the comparison. Figure 12 in Appendix further shows their interaction flow.
Table 3. Comparison of Three Intervention Methods
@@ 6.1. Baseline and MindShift-Simple Intervention Methods
Intent Report LLM-powered Persuasion Mental-States-Based Persuasion Strategies
@@
The table consists of four rows and four columns, including headers. The first column lists three ”Intervention Methods”: ”Baseline,” ”MindShift-Simple,” and ”MindShift.” The remaining columns describe each method’s ”Characteristics,” specifically ”Intent Report,” ”LLM-powered Persuasion,” and ”Mental-State-based Persuasion Strategies.” The first row is the primary header, and the second row serves as a subheader to specify characteristics. Row three details that ”Baseline” has just one characteristic: ”Intent Report.” Row four shows ”MindShift-Simple” has two characteristics: ”Intent Report” and ”LLM-powered Persuasion.” The final row indicates ”MindShift” incorporates all three characteristics.
@@ 6.2. Experiment Design
We adopted a within-subject design, with intervention techniques as independent variables (Baseline, MindShift-Simple, and MindShift). We designed a 5-week field experiment. To measure users’ everyday phone usage behavior, the first week is set as the Baseline stage, followed by two weeks of one MindShift version and another two weeks of the other version. We counter-balanced the order of MindShift-Simple and MindShift.
Our evaluation metrics include various aspects: (1) Intervention acceptance rate. We measure the percentage of times users accept the intervention and quit the blacklist app use when interventions are shown. (2) Intervention thumb-up rate. For MindShift-Simple and MindShift that show persuasion content, users can provide feedback by thumb-up or thumb-down (step 4 in Figure 5). (3) App usage behavior, which includes both app opening frequency and usage duration. (4) Subjective reports. At the beginning of the study and at the end of each intervention session, we distributed the Smartphone Addiction Scale (SAS) (Kwon et al., 2013) and the self-efficacy scale (Schwarzer and Jerusalem, 1995). In addition, at the end of the study, we also conducted a brief semi-structured interview to gather user experiences and feedback on our intervention techniques. These metrics cover both the objective and subjective measures of the interventions.
@@ 6.3. Participants
We sent out recruitment material on social media platforms. We included a screening survey aiming to identify potential participants who showed signs of smartphone addiction and the willing to reduce their smartphone use. Specifically, besides basic demographics, we included four questions selected from the SAS and self-efficacy questionnaires, questions about the willingness to reduce smartphone use, the extent of habitual phone use, future plans for the next five weeks, and a screenshot of phone usage time of the last week. We excluded users (1) without signs of smartphone addiction (SAS sub-score ¡ 15), or (2) unwilling to reduce smartphone use, or (3) less than 20 hours of weekly phone use, or (4) having special plans such as long-term travel in the next five weeks (which may shift their phone usage patterns).
We received a total of 42 responses. We recruited 31 participants after the screening process. 6 participants voluntarily dropped out during the study. For the remaining participants, we divided them into groups according to counterbalanced intervention orders. We conducted a Kruskal-Wallis test analysis to ensure that groups had no significant difference in the SAS scores and self-efficacy scores. In the end, 25 participants completed the entire study (females=13, males=12, age=22±2 years), including 17 undergraduates, 5 graduate students, and 3 professionals.
@@ 6.4. Experiment Procedure
After all the participants signed the consent form, we held a 20-minute onboarding session online to familiarize participants with the research process and introduce the Android application. We explained in detail the meanings of each selection in the intent report interfaces (Step 2 in Figure 5). After the meeting, participants filled out the first SAS and self-efficacy questionnaires. The app was then deployed for a 5-week field experiment.
Before users started using one of two versions of MindShift (MindShift and MindShift-Simple), we provided participants with a tutorial explaining the three mental states (i.e., boredom, stress, and inertia) that they need to report. To confirm their understanding, participants were asked to take a brief test on the understanding of mental states until they reached a score of 90%. This ensured that they had an accurate and consistent understanding of the mental states, which in turn improved the accuracy and reliability of the reported data. At the end of the experiment, participants received a compensation of \$50 for their time.
@@ 7. RESULTS
During the five-week study, we collected 50,815 minutes of restricted app usage duration, and 54,467 restricted app opening events (7539, 23,769, 21,994, and 835 for habitual use, instrumental use, relax, and quit). We conducted statistical tests on the quantitative data collected from the app and scale scores to measure differences. For qualitative data from exit interviews, we conducted thematic coding to extract key insights.
@@ 7.1. Intervention Acceptance Rate
The effectiveness of a persuasion strategy is directly measured by the rate of successful prevention of user engagement with the targeted application (i.e., intervention acceptance rate). We compared the overall acceptance rate (Sec. 7.1.1), the acceptance rate for generated persuasion content (Sec. 7.1.2), and thumb-up rate (Sec. 7.1.3). Moreover, we also conducted a detailed analysis of the acceptance rate across strategies (Sec. 7.1.4), mental states, and activities (Sec. 7.1.5).
@@ 7.1.1. Overall Acceptance Rate
(a) Overall Acceptance Rate (Session-based)
(b) Overall Acceptance Rate (Pop-up-based)
(Top) A bar plot of the session-based intervention acceptance rate of the three techniques. The X-axis is the three techniques. From the left to the right are Baseline, MindShift-Simple, and MindShift. The Y-axis is the acceptance rate(session-based). MindShift has the highest rate at 45.1%. MindShift-Simple has 38.5%, and Baseline has 12.7%. There are also two dashed lines on the bar of MindShift-Simple and MindShift. The upper line represents the first-round acceptance rate. MindShift has the highest rate at 38.7%, and MindShift-Simple has 36.5%. The lower line represents the intent-report-only acceptance rate. MindShift-Simple has the highest rate at 18.3%, and MindShift has 17.9%. (Bottom) A bar plot of the pop-up-based intervention acceptance rate of the three techniques. The X-axis is the three techniques. From the left to the right are Baseline, MindShift-Simple, and MindShift. The Y-axis is the acceptance rate(pop-up-based). MindShift has the highest rate at 35.2%. MindShift-Simple has 30.5%, and Baseline has 12.7%.
Figure 6. Overall Acceptance Rate
MindShift and MindShift-Simple increase the overall acceptance rate significantly and MindShift achieves best. We assess the overall acceptance rate using two methods. The first “session-based” rate means among total app visits (excluding instrumental uses and relaxation), how many times users quit during the intervention (including intent report and persuasion content  [^5] Users click the ”I am great and I will quit playing” in Step 1 or ”Quit using” in Step 3 in Fig 5). Figure 5(a) shows that MindShift (45.1±34.3%) achieves higher acceptance than MindShift-Simple (38.5±32.2%) and Baseline (12.7±17.4%). Significance is observed in a Friedman test ($\chi^{2}$(2) = 16.64, p ¡ .001). Three post hoc Wilcoxon signed-rank tests, corrected with Holm’s sequential Bonferroni procedure, indicate that MindShift vs. Baseline (V = 31, p ¡ .001) and MindShift-Simple vs. Baseline (V = 59, p ¡ .01) are significantly different, while MindShift vs. MindShift-Simple is not (V = 126, n.s.).
Considering that the Baseline doesn’t trigger subsequent interventions like the Mindshift and Mindshift-simple, we also compare the acceptance rates of the first round in particular. Both MindShift and MindShift-Simple initiate persuasion immediately after participants report their intents, so we include this initial persuasion in calculating their first-round acceptance rates. Additionally, we analyze the acceptance rate of the intent report to distinguish its effectiveness among the three intervention techniques. As shown in the upper dashed lines in Figure 5(a), the first-round acceptance rates for MindShift (38.7%±24.9%) and MindShift-Simple (36.2%±25.8%) are still significantly higher than the Baseline (12.7%±17.4%,$\chi^{2}$(2) = 15.56, p ¡ .001). Post hoc tests show significant differences in MindShift vs. Baseline (V = 43, p ¡ .01), and MindShift-Simple vs. Baseline (V = 32, p ¡ .001), but not in MindShift vs. MindShift-Simple (V = 134, n.s.). The lower dashed lines in Figure 5(a) indicate that the acceptance rates when considering only reporting intent are still higher for MindShift (17.9%±17.5%) and MindShift-Simple (18.3%±18.2%) compared to the Baseline (12.7%±17.4%). However, a Friedman test ($\chi^{2}$(2) = 5.59, p ¡ .1) does not show significance, suggesting that intent report has no difference among the three intervention techniques.
Following the session-based method, we also investigate the pop-up-based acceptance rate, which equals the total number of quit times divided by the total number of intervention pop-ups (i.e., each round is counted as a pop-up). Figure 5(b) shows the comparison, MindShift (35.2±24.1%) still has a higher acceptance than MindShift-Simple (30.5±23.6%) and Baseline (12.7±17.4%, $\chi^{2}$(2) = 13.69, p ¡ .01). Post hoc tests indicate significance for MindShift vs. Baseline (V = 52, p ¡ .01) and MindShift-Simple vs. Baseline (V = 38, p ¡ .001), but not for MindShift vs. MindShift-Simple (V = 129, n.s.). Subsequent analyses are all based on pop-ups.
@@ 7.1.2. Acceptance Rate for Generated Persuasion Content
(Top) A bar plot of the persuasion acceptance rate of two versions of MindShift. The X-axis is the two techniques. From left to right are MindShift-Simple, and MindShift. The Y-axis is the persuasion acceptance rate. MindShift has the highest rate of 20.1%. MindShift-Simple has 12.0%. (Bottom) A bar plot of the persuasion acceptance rate of two versions of MindShift grouped by 4 rounds. The X-axis is the round. From the left to the right are round1, round2, round3, round4. And each round has two bars, the left is MindShift-Simple while the right is MindShift. The Y-axis is the persuasion acceptance rate. MindShift has a higher acceptance rate in every round.
@@
MindShift has a significantly higher persuasion acceptance rate than MindShift-Simple. The overall acceptance rate includes two parts: exiting when reporting intent and exiting after seeing the persuasion content. To narrow down the comparison between MindShift and MindShift-Simple, we exclude the intent report stage and focus on the acceptance rate during the persuasion stage, as they differ only in the persuasive content. As shown in Figure 6(a), MindShift achieves higher persuasion acceptance (20.1±20.2%) than MindShift-Simple (12.0±15.0%) and a paired-samples t-test shows that MindShift was statistically significantly higher (t = -2.21, p¡0.05).
Moreover, as we introduce in Sec. 4.3, every persuasion intervention could consist of 1 to 4 rounds, depending on which mental state participants are in and which stage participants leave the app. Therefore, we further compare the persuasion acceptance rates between MindShift and MindShift-Simple across different persuasion rounds. The results show that MindShift outperforms MindShift-Simple at each round  [^6] Round 1 here only contains persuasion and excludes the intent report.($\Delta_{\textit{Round1}}$=6%, $\Delta_{\textit{Round2}}$=14.7%, $\Delta_{\textit{Round3}}$=10.2%, $\Delta_{\textit{Round4}}$=11%) as indicated in Figure 6(b). We conduct a paired-samples t-test between the two intervention techniques in each round. Results indicate that, except for the first round, rounds 2 (p¡0.05), 3 (p¡0.001), and 4 (p¡0.001) all exhibit that MindShift achieves significantly higher acceptance rate. This trend suggests that as the number of interventions increases, MindShift’s advantage becomes more pronounced, highlighting MindShift’s robustness and effectiveness in maintaining high acceptance rates.